OpenStudy (anonymous):

√(x+2) + 4=x Show work

OpenStudy (anonymous):

(x+2)^(1/2) = x-4 square both sides x+2 = x^2 -8x + 16 x^2 -9x + 14 = 0 therefore on solving x = 2 or, 7

OpenStudy (anonymous):

Careful !!! All answers to radical equations MUST BE CHECKED. x=7 checks OK because (7+2)^(1/2) + 4 =7 9^(1/2)+4=7 3+4=7 7=7 YES But x=2 must be REJECTED because it does not check. (2+2)^(1/2)+4=2 4^(1/2)+4=2 2+4=2 6=2 No Way! Reject x=2 visit http://www.algebraINAflash.com